Archived Service Status (2006)
Wednesday July 24, 2024 - 21:40:02
Orpheus Internet Services Ltd

In case of server failure, this page is mirrored at

Archived Operational Issues (2007)

Your IP address is currently (full info)

Return to current status

Normal service resumed
Thu 17th Jan 2008
pvNormal service should be resumed after this afternoon's server outage. Sorry for any inconvenience.
WebMail fixed
Thu 15th Nov 2007
pvWebMail is now fixed and all services running normally again. Sorry for any inconvenience.
WebMail problem
Thu 15th Nov 2007
pvWebMail is temporarily unavailable. We're currently investigating the problem. Some hosted websites may also be affected.
WebMail fixed
Tue 9th Oct 2007
pvWebMail is now available. Apologies for any temporary inconvenience.
WebMail unavailable
Tue 9th Oct 2007
pvWebMail is currently unavailable. We are investigating the problem. Other email services (POP3/IMAP) are working normally.
Dialup issues
Sun 16th Sep 2007
pvSome people are reporting problems connecting to the 0845 dialup number. We've reported this as a fault to telecom engineers and await a status report. It seems localised to some regional exchanges.
Usenet upgraded
Sat 25th Aug 2007
pvI've upgraded the news server to be more efficient and more flexible too. I'm still carrying out some tests, but everything should be working normally now.
Usenet maintenance
Sat 25th Aug 2007
pvWe're currently doing some routing maintenance on so our Usenet feed may be intermittent this morning.
WebMail restored
Tue 14th Aug 2007
pvWebMail has now been restored
News restored
Tue 14th Aug 2007
02.45pm should be working again for local orpheus.* newsgroups
SMTP restored - WebMail temporarily unavailable
Tue 14th Aug 2007
pvSMTP is restored, so mail is working normally. However, due to server failure, WebMail and news are currently unavailable. Normal service will be resumed as soon as possible.
SMTP problems
Tue 14th Aug 2007
pvThere is currently an issue with the SMTP server, we're currently investigating the problem.
WebMail restored
Wed 8th Aug 2007
pvThe WebMail system is restored and working normally.
WebMail temporarily unavailable
Wed 8th Aug 2007
pvThe WebMail system is currently being intermittent. We are investigating the problem and normal service will be resumed as soon as possible.
Usenet temporarily suspended
Wed 25th Jul 2007
pvOur Usenet feed is currently over quota, so external newsgroups may be unavailable/intermittent. In the meantime, orpheus.* newsgroups are working normally.
Reduced capacity on Usenet
Wed 18th Jul 2007
pvDue to reaching 99% of our Usenet quota, I've temporarily disabled the uk.* hierarchy in an attempt to make our remaining quota last a bit longer. Normal service will be resumed around 15th August.
See our Usenet page for more info.
Usenet feed restored
Mon 16th Jul 2007
pvOur newsfeed has been restored, but unfortunately, due to excessive use I've temporarily suspended alt.* newsgroups, which was draining our entire Usenet resources.
Usenet temporarily suspended
Tue 10th Jul 2007
pvUnfortunately, our Usenet feed is still unavailable. It will return on Sunday 15th July. Sorry for any inconvenience.
Usenet temporarily suspended
Sun 1st Jul 2007
pvOur Usenet feed is currently over quota due to misuse. I'm currently investigating the cause, and endeavouring to find a prompt solution.
Usenet restored
Sat 23rd Jun 2007
pvI've now restored the news feed, so all services should now be operating normally.
Usenet unavailable
Sat 23rd Jun 2007
pvNews is currently unavailable.
Routine maintenance completed
Thur 21st Jun 2007
pvMaintenance is now complete and all services are working normally. We'd like to apoligise for any temporary inconvenience during our server upgrade.
Routine maintenance in progress
Thu 21st Jun 2007
pvOur planned maintenance has just got underway, meaning that Orac (POP mail, some websites) will be unavailable for approx 2 hours. Mail will be queued on the backup server, so no messages will be lost.
DNS issues resolved
Wed 20th Jun 2007
pvThe earlier DNS issues have been resolved and maintenance has been completed. All systems now working normally. Sorry for any inconvenience.
DNS issues
Wed 20th Jun 2007
pvSome hosted websites are currently failing to resolve correctly, due to DNS maintenance. We're investigating the problem and normal service will be resumed as soon as possible.
Wed 20th Jun 2007
01.30pm - 03.30pm
pvWe will be doing some brief maintenance today, so please check the maintenance page.
SMTP unavailable
Wed 6th Jun 2007
pvThe SMTP server is currently being updated. Normal service will be resumed ASAP.
Normal service resumed at 5:00pm
Usenet unavailable
Wed 6th Jun 2007
1.30pm - 02:12pm
pvUsenet is currently unavailable whilst we upgrade our news server.
Normal service resumed at 2:12pm
Possible broadband connection problems
Thu 10th May 2007
1.30pm - 2.00pm
pvI've just received notification that there is, and I quote, "Central pipe emergency maintenance work following an unplanned BT incident". I don't know what the 'incident' is, and it's not specifically aimed at Orpheus customers, but ADSL users in general, so I thought I'd report it here in case people are experiencing any unusual connection problems today.
Essential Maintenance
Thu 22nd Mar 2007
00.01am - 04:00am
pvBT are carrying out some essential maintenance work which may result in approx 30% of ADSL circuits being down for approx 20 minutes during the alloted period. Most routers should automatically reconnect, but if you encounter problems, try power-cycling the router or phoning the support line.
WebMail fixed
Wed 28th Feb 2007
pvWebMail is now restored. However some of the user preferences may have been reset. Please check your settings when you login. If you can't connect, try
WebMail problem
Wed 28th Feb 2007
pvWebMail is currently unavailable. I'm investigating the problem....
POP restored
Mon 12th Feb 2007
pvAll services have now been restored. Many apologies for the temporary unavailablity over night.
POP failure
Mon 12th Feb 2007
pvThe POP server seems to have temporarily failer. Engineers are currently working to fix the problem. SMTP and other services are working normally, although some websites may be unavailable.
Technical Helpdesk: Email: 
24 hour emergency:
0845 120 5559
07913 398 001 (voice or sms)

Download TeamViewer Download TeamViewer
Valid HTML 4.01! Valid CSS!